12. Write each radical expression in simplified radical form. sqrt2/sqrt18
60. Simplify. (3sqrt3)^4

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

12. Write each radical expression in simplified radical form.
sqrt2/sqrt18 = sqrt[1/9) = 1/3
=====================================

60. Simplify. (3sqrt3)^4 = 81*9 = 729
===========================================
Cheers,
Stan H.

Answer by edjones(8007)

sqrt(2)/sqrt(18)
=sqrt(2)/sqrt(9*2)
=sqrt(2)/3sqrt(2)
=1/3
.
Ed